which an equation of the line that passes though the point (5,2) and (10,-3)

Found 3 solutions by josgarithmetic, math_tutor2020, ikleyn:
Answer by josgarithmetic(39617)

typical, common,
look for other already solved examples.

Answer by math_tutor2020(3817)


Let's find the slope first.
(x1,y1) = (5,2) and (x2,y2) = (10,-3)
m = slope
m = rise/run
m = (change in y)/(change in x)
m = (y2 - y1)/(x2 - x1)
m = (-3 - 2)/(10 - 5)
m = (-5)/(5)
m = -1
The slope is -1

Now let's find the y intercept
I'll use the slope we just found and the first point (x1,y1) mentioned.
Use point-slope form
y-y1 = m(x-x1)
y-2 = -1(x-5)
y-2 = -1*x-1*(-5)
y-2 = -x+5
y = -x+5+2
y = -x+7

This equation is in y = mx+b form
m = -1 = slope
b = 7 = y intercept

Answer: y = -x+7
To verify this answer, plug in each coordinate into the equation. Then simplify. You should get the same thing on both sides.

Answer by ikleyn(52781)

.
which an equation of the line that passes though the point (5,2) and (10,-3).
~~~~~~~~~~~~~~~~


It is very elementary and basic skill that every student must develop to the level to solve such problems
on his  (or her)  own without asking for assistance from outside.


To help you to develop such skills,  special lesson was created in this site
    - Equation for a straight line in a coordinate plane passing through two given points


Read it and then solve your problem.


From this lesson,  learn the subject once and for all.


////////////////


At the end of this lesson,  you will find references  (the links)  to other lessons,
that cover the entire subject.

If you want to learn the entire subject,  do not miss these lessons - - - look into them.


Then you will see sunrise  ( ! )